# Audio-Visual Compound Expression Recognition Method based on Late Modality Fusion and Rule-based Decision

### v1

@@ Abstract
This paper presents the results of the SUN team for the Compound Expressions (CE)Recognition Challenge of the 6th ABAW Competition. We propose a novel audio-visual method for compound expression recognition. Our method relies on emotion recognition models that fuse modalities at the emotion probability level, while decisions regarding the prediction of compound expressions are based on predefined rules. Notably, our method does not use any training data specific to the target task. The method is evaluated in multi-corpus training and cross-corpus validation setups. Our findings from the challenge demonstrate that the proposed method can potentially form a basis for development of intelligent tools for annotating audio-visual data in the context of human’s basic and compound emotions. The source code is publicly available [^1] https://c-expr-db.github.io/AVCER/.
@@ 1 Introduction
The Compound Expression Recognition (CER) as a part of affective computing is a crucial task in intelligent human-computer interaction and multimodal user interfaces. It entails the automated identification of compound emotional states in individuals, which may include combinations of two or more basic emotions such as: Fear, Happiness, Sadness, Anger, Surprise, and Disgust.
Over the last two decades, research efforts in the field of automatic expression analysis have predominantly focused on identifying six basic emotions [30, 9]. However, these methods fail to fully capture a complexity of everyday emotional expressions. Individuals often exhibit Compound Expressionss (CEs), such as Fearfully Surprised, Happily Surprised, Sadly Surprised, Disgustedly Surprised, Angrily Surprised, Sadly Fearful, Sadly Angry, which are combinations of basic emotions. These CEs underscore the necessity for more comprehensive models capable of capturing the subtleties inherent in human’s emotional expressions.
@@
In this paper, we present a novel method for audio-visual CER as a part of the 6th ABAW CE Recognition Challenge. Our method does not utilize the CER challenge data as training data; rather, it includes models trained for basic emotion recognition. Decisions regarding predicted CEs are determined by the proposed rule-based method. It enables us to address the problem of insufficient publicly available data with CEs. Moreover, CEs comprise various pair combinations of basic emotions, rendering the proposed method particularly valuable to the research community. The versatility of the proposed method allows using as a software tool for annotating new emotionally coloured data.
@@ 2.3 Modality Fusion
where $v_{1}\in V$. $W$ and $V$ are weights that remain consistent across all test samples. This hierarchical weighting enhances performance measures for both basic emotion recognition and CE recognition by considering the contribution of each model. The final probability vector is then utilized for CER.
@@ 2.4 Rule-based Decision-Making Method
4 Model ID 1 & 2 (hierarchical weighing) 42.27 46.19 43.79 44.84 XX XX
@@
8 Model ID 1 & 2 & 5 (hierarchical weighing) 42.77 49.98 43.09 43.55 XX
9 Model ID 1 & 2 & 6 (Dirichlet-based weighing) 46.79 51.79 44.81 45.66 XX XX
@@ 3.2 Experimental Results
The experimental results are presented in Table 2. Depending on the corpus, the visual models show different performance; the dynamic model outperforms the static model for AffWild2, and vice versa for AFEW. When the two visual models are combined, the hierarchical weighting fusion outperforms the Dirichlet-based weighting. Therefore, in the first submission to the CE Recognition Challenge, we use Model ID 4 in two versions: with and without the Rule 2 to predict CE. We then evaluate the performance of the audio models on two corpora, AffWild2 and AFEW. It is impossible to say which model is more efficient. However, since these models show lower performance than the video models, we do not test only the audio models on C-EXPR-DB. Finally, we fuse the audio and video models by comparing various weighting fusions. The results show that the hierarchical weighting fusion outperforms the Dirichlet-based fusion. However, for comparison with previous models tested on C-EXPR-DB, we submit Model ID 8 with the Rule 2 for CE prediction. Additionally, we evaluate Model ID 9 based on Dirichlet fusion with both rules.
@@
Therefore, we have developed a method comprising three models, each assigned responsibility for predicting its respective class during CER. The proposed method can serve as a tool for annotating both basic emotions and CEs.

### v2

@@ Abstract
This paper presents the results of the SUN team for the Compound Expression (CE)Recognition Challenge of the 6th ABAW Competition. We propose a novel audio-visual method for compound expression recognition. Our method relies on emotion recognition models that fuse modalities at the emotion probability level, while decisions regarding the prediction of compound expressions are based on predefined rules. Notably, our method does not use any training data specific to the target task. Thus, the problem is a zero-shot classification task. The method is evaluated in multi-corpus training and cross-corpus validation setups. Using our proposed method is achieved an F1-score value equals to 22.01% on the C-EXPR-DB test subset. Our findings from the challenge demonstrate that the proposed method can potentially form a basis for developing intelligent tools for annotating audio-visual data in the context of human’s basic and compound emotions.
@@ 1 Introduction
The Compound Expression Recognition (CER) as a part of affective computing is a novel task in intelligent human-computer interaction and multimodal user interfaces. It entails the automated identification of compound emotional states in individuals, which may include combinations of two or more basic emotions such as: Fear, Happiness, Sadness, Anger, Surprise, and Disgust.
Over the last two decades, research efforts in the field of automatic expression analysis have predominantly focused on identifying six basic emotions [30, 9]. However, these methods fail to fully capture the complexity of everyday emotional expressions. Individuals often exhibit Compound Expressions (CEs), such as Fearfully Surprised, Happily Surprised, Sadly Surprised, Disgustedly Surprised, Angrily Surprised, Sadly Fearful, Sadly Angry, which are combinations of basic emotions. This CEs underscores the necessity for more comprehensive models capable of capturing the subtleties inherent in human’s emotional expressions.
@@
In this paper, we present a novel method for audio-visual CER as a part of the 6th ABAW CE Recognition Challenge. Our method does not utilize the CER challenge data as training data; rather, it includes models trained for basic emotion recognition. Decisions regarding predicted CEs are determined by the proposed rule-based method. It enables us to address the problem of insufficient publicly available data with CEs. Moreover, CEs comprise various pair combinations of basic emotions, rendering the proposed method particularly valuable to the research community. The versatility of the proposed method allows for use as a software tool for annotating new emotionally colored data.
@@ 2.3 Modality Fusion
where $v_{1}\in V$. $W$ and $V$ are weights that remain consistent across all test samples. This hierarchical weighting enhances performance measures for basic emotion recognition by considering the contribution of each model. The final probability vector is then utilized for CER.
@@ 2.4 Rule-based Decision-Making Method
4 Model ID 1 & 2 (hierarchical weighing) 42.27 46.19 43.79 44.84 19.44 20.98
@@
8 Model ID 1 & 2 & 5 (hierarchical weighing) 42.77 49.98 43.09 43.55 17.56
9 Model ID 1 & 2 & 6 (Dirichlet-based weighing) 46.79 51.79 44.81 45.66 21.14 22.01
@@ 3.2 Experimental Results
The experimental results are presented in Table 2. Depending on the corpus, the visual models show different performance; the dynamic model outperforms the static model for AffWild2, and vice versa for AFEW. When the two visual models are combined, the hierarchical weighting fusion outperforms the Dirichlet-based weighting. Therefore, in the first submission to the CE Recognition Challenge, we use Model ID 4 in two versions: with and without the Rule 2 to predict CE. Using Rule 2 reduces the performance measure of CER by 1.54% compared to using the rule-free method. This suggests that increasing the importance of the less represented emotions in CE pairs does not improve CER performance.
We then evaluate the performance of the audio models on two corpora, AffWild2 and AFEW. It is impossible to say which model is more efficient. However, since these models show lower performance than the video models, we do not test only the audio models on C-EXPR-DB. Finally, we fuse the audio and video models by comparing various weighting fusions. The results show that the hierarchical weighting fusion outperforms the Dirichlet-based fusion. However, for comparison with previous models tested on C-EXPR-DB, we submit Model ID 8 with the Rule 2 for CE prediction. Additionally, we evaluate Model ID 9 based on Dirichlet fusion with both rules. Combining the visual models with the audio model results in a 2.57% improvement in F1. However, employing hierarchical weighting for model fusion shows a performance decrease of 4.45% compared to Dirichlet-based weighting. Additionally, Rule 1 exhibits a 0.87% lower performance than Rule 2.
@@
Therefore, we have developed a method comprising three models, each assigned responsibility for predicting its respective class during CER.
@@ 4 Conclusions
5 Acknowledgements
This study was supported by the Russian Science Foundation (project No. 22-11-00321).